OpenStudy (anonymous):

What is the vertex of a parabola in the form of the equation y=a(x – h)2 + k?

OpenStudy (precal):

(h, k) is your vertex

OpenStudy (precal):

if a is + then it is a min if a is - then it is a max

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Solve. x2 + 8x = 24 ? I AM NOT GOOD WITH MATH--Can you halp me?

OpenStudy (anonymous):

me the same precal

OpenStudy (precal):

first move 24 to the other side x^2 + 8x - 24=0

OpenStudy (precal):

second you can use the vertex formula x=-b/2a

OpenStudy (anonymous):

I see ... what is next ?

OpenStudy (precal):

 we now know the x value of your vertex by the way it is a min